OpenStudy (simplyqveen038):

What is the equation of the line that passes through the points (−2, 1) and (1, 10)? A: 3x-y=-7 B:3x-y=-5 C:3x-y=5 D:x+3y=-5

satellite73 (satellite73):

you know how to find the slope?

OpenStudy (simplyqveen038):

yes m= y2-y1 / x2=x1

satellite73 (satellite73):

ok you need that first

OpenStudy (simplyqveen038):

i did that and i got 3

satellite73 (satellite73):

ok, next comes point slope formula

OpenStudy (simplyqveen038):

i thought it was letter A but its negative not positive

satellite73 (satellite73):

let's pick \((1,10)\) so we start with \[y-y_1=m(x-x_1)\] or \[y-10=3(x-1)\]

satellite73 (satellite73):

multiply out get \[y-10=3x-3\] subtract \(3x\) from both sides i guess

OpenStudy (simplyqveen038):

i thought you would have to add 10 to 3?

satellite73 (satellite73):

sure you get \]-3x+y=7\]

satellite73 (satellite73):

oops i meant \[-3x+y=7\]

OpenStudy (simplyqveen038):

but thats not one of the choices :(

satellite73 (satellite73):

course that is not one of your answer choices multiply both sides by \(-1\) get \[3x-7=-7\]

OpenStudy (simplyqveen038):

im confused

OpenStudy (simplyqveen038):

hello